OpenStudy (anonymous):

Write the expression as a single logarithm. 5log q + 8log t b b

OpenStudy (anonymous):

bring the numbers out the front , and but them as powers

OpenStudy (anonymous):

\[n \log_{b} x = \log_{b} x^n \]

OpenStudy (anonymous):

\[= \log_{b} q^5 + \log_{b} t^8\]

OpenStudy (anonymous):

Also \[\log(a) + \log(b ) = \log(ab) \]

OpenStudy (anonymous):

so = \[= \log_{b} (q^5 t^8)\]